### v2

@@ 4.2 Methods
We then consider five adaptation methods in our study, ranging from linear probing methods to partial fine-tuning and full fine-tuning methods:
Linear: A common approach [27, 33] consists in freezing the pretrained backbone and adding a segmentation head, called decoder, consisting of a batch normalization layer followed by a convolutional 1x1 layer. Figure 2 illustrates the overall pipeline for this linear adaptation method.
Multilayer: It has been shown in [27] and [38] that using a concatenation of multiple layers from the pretrained model can boost performance in downstream tasks. We thus propose to use a linear method on the concatenation of the last 4 blocks of the pretrained model.
Singular Value Fine-tuning [34] (SVF): The SVF method decomposes parameters into three matrices using Singular Value Decomposition (SVD) and fine-tunes only the singular values diagonal matrix while keeping the rest frozen. In [34], this approach demonstrated promising results in few-shot semantic segmentation tasks, showcasing its ability to specialize the pretrained model without overfitting. In our study, we initially apply the linear method and then proceed to fine-tune the singular values of the encoder alongside the decoder parameters.
Low-Rank Adaptation [14] (LoRA): In this approach, we maintain the pretrained backbone’s frozen state and introduce low-rank trainable adapters, as described in [14]. Similarly to the SVF method, LoRA enables tuning only a limited number of parameters. Our procedure begins with training the decoder using the linear method and then introduces LoRA adapters to the encoder. Subsequently, we fine-tune both the adapters and decoder parameters.
Fine-tuning: The fine-tuning method, a straightforward yet effective approach, involves fine-tuning the entire pretrained model. This method has demonstrated state-of-the-art performance in few-shot classification [21]. However, it is susceptible to overfitting due to the limited number of training samples compared to the typically extensive parameter pool in the considered encoders.
@@ 5.1 Training Procedure
To ensure the reproducibility of our results, we conducted all experiments with three fixed random seeds and computed the mean Intersection-over-Union (mIoU) across these runs. For preprocessing, we followed the approach outlined in [10]. This involved random horizontal flipping, random scaling of the shorter side within the range of [400, 1600] while preserving the aspect ratio, and random cropping to a size of 1024x1024. We applied a subset of RandAug [4] operations, including auto contrast, equalize, rotate, color, contrast, brightness, and sharpness. Subsequently, all images were resized to 1024x1024 resolution, except for CLIP and MAE, which accept input sizes of 224x224. We acknowledge that the resolution variation, depending on the considered model, may introduce some analysis challenges but was an unavoidable aspect of our setup. For the learning rate, we employed the poly learning rate scheduler with a power of 0.9 for the linear and multilayer methods. Specifically, we used a learning rate of 0.2 for Cityscapes, 0.05 for COCO, and 0.001 for PPD. For fine-tuning methods (SVF, LoRA, and fine-tune), we did a grid search across models and datasets with values including $10^{-2}$, $10^{-3}$, $10^{-4}$, $10^{-5}$, and $10^{-6}$ to determine the one yielding the best average performance. Linear and multilayer methods were trained with a batch size of 4. For the Fine-tuning, SVF, and LoRA methods, we initiated the training with the linear method and subsequently performed fine-tuning with the reduced batch size of 2, addressing memory limitations. Training duration extended to 200 epochs for Cityscapes and PPD, and 100 epochs for COCO. The query set consisted of the standard validation sets for Cityscapes and COCO. For PPD, given the absence of a conventional split, we performed a fixed 50/50 random split into training and validation sets. Our models were trained on a combination of NVIDIA A6000 and NVIDIA RTX3090 GPUs. Runtime varied across datasets and models, ranging from approximately 30 minutes per run for MAE, CLIP, and ResNet on Cityscapes to approximately 4 hours per run for DINO V2 on COCO. For the size of the models, we chose the base version of the vision transformers, leading to a 89M parameters model for SAM and 86M for DINO V2, CLIP and MAE. For the FCN we chose a ResNet50 as an encoder, leading to a 23M parameters model.
@@ 5.2 What Works The Best?
First, we aim to identify the optimal combination of pretrained models and adaptation methods based on mIoU. Table 1 presents the average performance on the three considered datasets for all possible combinations of pretrained models and adaptation methods, along with their standard deviation, in the context of 1-shot problems. Several conclusions can be drawn:
  • DINO V2 consistently outperforms all other models across various settings. This superiority is especially notable on Cityscapes and COCO, where DINO V2 surpasses other models by a significant margin (see Appendix Table 9 for the detailed results). It is worth noting that Cityscapes is included in the training images of DINO V2, although it only had access to the images and not their associated segmentation masks.
  • SAM encoder exhibits strong performance on average. It outperforms CLIP, MAE, and ResNet on both Cityscapes and PPD. However, it yields comparatively poor results on COCO (refer to Appendix Table 9). This difference is likely attributed to our decision to retain the SAM image encoder while discarding the original mask decoder.
  • The multilayer method outperforms the linear approach for most models, indicating that including earlier feature maps, which are more locally focused [29, 7], provides the decoder with greater segmentation flexibility. Furthermore, it exhibits a lower standard deviation, indicating greater stability across the different runs of our proposed benchmark.
  • Despite variations in the number of trainable parameters (less than 1% for LoRA, around 2% for SVF, and 100% for finetune), the finetuning methods (SVF, LoRA, and Full Fine-tuning) yield comparable results between them.
Table 1: mIoU averaged over our three considered datasets in 1-shot few-shot semantic segmentation. Bold numbers correspond to the best performance across feature extractors and underline numbers to the best performance across adaptation methods.
@@
We then vary the number of shots to 2, 5 and 10. Table 2 shows the consistency of the results with 1-shot, where:
  • DINO V2 outperforms other models by a big margin.
  • SAM still yields strong performance on average, but performs poorly on COCO (refer to appendix Tables 101112).
  • The multilayer method outperforms the linear approach for most models.
  • Fine-tuning methods yield comparable results.
In addition, we can see that:
  • DINO V2 scales better than other models with more shots.
  • Fine-tuning scales better than LoRA and SVF and gives better results on average on the 10-shot setting, this is due to the fact that we suffer less from over-fitting when having more shots.
Table 2: mIoU averaged over our three considered datasets in $k$-shot semantic segmentation for $k\in\{2,5,10\}$. Bold numbers correspond to the best performance across extractors and underline numbers to the best performance across adaptation methods.
Linear Multilayer SVF LoRA Fine-tuning

### v3

@@ 5.2 Methods
We then consider five adaptation methods in our study, ranging from linear probing methods to partial fine-tuning and full fine-tuning methods  [^5] We do not consider meta-learning methods in our study, as these approaches rely on base classes to facilitate rapid adaptation to new classes, such as in episodic training. In contrast, our benchmark directly begins training on the support set.:
Linear: A common approach (Oquab et al., 2023; Strudel et al., 2021) consists in freezing the pretrained backbone and adding a segmentation head, called decoder, consisting of a batch normalization layer followed by a convolutional 1x1 layer. Figure 2 illustrates the overall pipeline for this linear adaptation method.
Multilayer: It has been shown in (Oquab et al., 2023) and (Xie et al., 2021) that using a concatenation of multiple layers from the pretrained model can boost performance in downstream tasks. We thus propose to use a linear method on the concatenation of the last 4 blocks of the pretrained model.
Singular Value Fine-tuning (Sun et al., 2022) (SVF): The SVF method decomposes parameters into three matrices using Singular Value Decomposition (SVD) and fine-tunes only the singular values diagonal matrix while keeping the rest frozen. In (Sun et al., 2022), this approach demonstrated promising results in few-shot semantic segmentation tasks, showcasing its ability to specialize the pretrained model without overfitting. In our study, we initially apply the linear method and then proceed to fine-tune the singular values of the encoder alongside the decoder parameters.
Low-Rank Adaptation (Hu et al., 2022) (LoRA): In this approach, we maintain the pretrained backbone’s frozen state and introduce low-rank trainable adapters, as described in (Hu et al., 2022). Similarly to the SVF method, LoRA enables tuning only a limited number of parameters. Our procedure begins with training the decoder using the linear method and then introduces LoRA adapters to the encoder. Subsequently, we fine-tune both the adapters and decoder parameters.
Fine-tuning: The fine-tuning method, a straightforward yet effective approach, involves fine-tuning the entire pretrained model. This method has demonstrated state-of-the-art performance in few-shot classification (Luo et al., 2023). However, it is susceptible to overfitting due to the limited number of training samples compared to the typically extensive parameter pool in the considered encoders.
@@ 6.1 Training Procedure
To ensure the reproducibility of our results, we conducted all experiments with three fixed random seeds and computed the mean intersection-over-union (mIoU) across these runs. For the results averaged across the datasets, we averaged the results per run corresponding to a seed and thus having the same sampled support set, and then averaged these three means to compute their mean and standard deviation. For preprocessing, we followed the approach outlined in (Gao, 2023). This involved random horizontal flipping, random scaling of the shorter side within the range of [400, 1600] while preserving the aspect ratio, and random cropping to a size of 1024x1024. We applied a subset of RandAug (Cubuk et al., 2020) operations, including auto contrast, equalize, rotate, color, contrast, brightness, and sharpness. Subsequently, all images were resized to 1024x1024 resolution, except for CLIP and MAE, which accept input sizes of 224x224. We acknowledge that the resolution variation, depending on the considered model, may introduce some analysis challenges but was an unavoidable aspect of our setup. For the learning rate, we employed the Polynomial learning rate scheduler with a power of 0.9 for the linear and multilayer methods. Specifically, we used a learning rate of 0.2 for Cityscapes, 0.05 for COCO, and 0.001 for PPD. For fine-tuning methods (SVF, LoRA, and fine-tune), we did a grid search across models and datasets with values including $10^{-2}$, $10^{-3}$, $10^{-4}$, $10^{-5}$, and $10^{-6}$ to determine the one yielding the best average performance for 1-shot. Once identified, we applied this optimal learning rate to experiments involving more shots, primarily due to computational constraints. Linear and multilayer methods were trained with a batch size of 4. For the Fine-tuning, SVF, and LoRA methods, the training phase consists of two stages. First, we train the decoder using the linear method. Then, we perform fine-tuning on both the decoder and the adapters (or all parameters in the case of fine-tuning) with a reduced batch size of 2 to address memory constraints. Training duration extended to 200 epochs for Cityscapes and PPD, and 100 epochs for COCO. The query set consisted of the standard validation sets for Cityscapes and COCO. For PPD, given the absence of a conventional split, we performed a fixed 50/50 random split into training and validation sets. Our models were trained on a combination of NVIDIA A6000, NVIDIA RTX3090 and NVIDIA A100 GPUs. Runtime varied across datasets and models, ranging from approximately 30 minutes per run for MAE, CLIP, and ResNet on Cityscapes to approximately 4 hours per run for DINOv2 on COCO. For the size of the models, we chose the base version of the vision transformers, leading to a 89M parameters model for SAM and 86M for DINOv2, CLIP and MAE. For the FCN we chose a ResNet50 as an encoder, leading to a 23M parameters model.
@@ 6.2 What Works The Best?
First, we aim to identify the optimal combination of pretrained models and adaptation methods based on mIoU. Table 1 presents the average performance on the three considered datasets for all combinations of pretrained models and adaptation methods, along with their standard deviation over 3 runs (6.1), in the context of $\{1,2,5,10\}$-shots. Several conclusions can be drawn:
  • DINOv2 consistently outperforms all other models across various settings. This superiority is especially notable on Cityscapes and COCO, where DINOv2 surpasses other models by a significant margin (see Appendix Tables 3, 4, 5, 6 for the detailed results). Surprisingly, it outperforms supervised models pre-trained on other semantic segmentation datasets, showing the potential superiority of VFM adaptation over cross-domain generalization. However, since Cityscapes is included in DINOv2’s pretraining data, albeit without segmentation masks, some degree of data leakage may be present. As stated in the introduction, our primary objective is to assess the adaptation capabilities of off-the-shelf VFM rather than their generalization abilities. Recent studies (Udandarao et al., 2024) have begun exploring this issue, emphasizing the need for further research into the influence of pretraining data distribution on downstream tasks.
  • SAM encoder exhibits strong performance on average. It outperforms CLIP, MAE, and ResNet on both Cityscapes and PPD. However, it yields comparatively poor results on COCO (refer to Appendix Tables 3, 4, 5, 6). This difference is likely attributed to our decision to retain the SAM image encoder while discarding the original mask decoder A.3.
  • The multilayer method outperforms the linear approach for most models, indicating that including earlier feature maps, which are more locally focused (Raghu et al., 2021; Dosovitskiy et al., 2021), provides the decoder with greater segmentation flexibility. Furthermore, it exhibits a lower standard deviation, indicating greater stability across the different runs of our proposed benchmark.
  • Despite variations in the number of trainable parameters (less than 1% for LoRA, around 2% for SVF, and 100% for finetune), the finetuning methods (SVF, LoRA, and Full Fine-tuning) yield comparable results between them highlighting the importance of using efficient methods.
  • Fine-tuning scales better than LoRA and SVF and gives better results on average with 10 shots, this is due to the fact that we suffer less from over-fitting when having more shots.
Table 1: mIoU averaged over our three considered datasets in $k$-shot semantic segmentation for $k\in\{1,2,5,10\}$. Bold numbers correspond to the best performance across extractors and underline numbers to the best performance across adaptation methods.
@@
SegFormer 45.60$\pm$02.40 48.96$\pm$01.90 50.92$\pm$02.41 49.93$\pm$01.23 51.16$\pm$01.83